A firm produces two commodities, ×, and y. The demand functions are

P, = 900 - 2x - 2 y

and

P, = 1400 - 2x -4y

respectively, where P, is the price of commodity x and P, is the price of commodity y. The costs

are given by

C, = 7000 + 100x + x2

and

C, = 10000 + 6 y*

a) Show that the firm's profit function is given by

7 (x, y) = -3x? - 10y? 4xy + 800x +1400y -17000

b) Suppose the firm is required to produce a total of exactly 60 units. Find the values of x

and y that maximize profits.


1
Expert's answer
2022-05-23T23:42:52-0400

a)


"\pi(x, y)=x(900 - 2x - 2 y)-(7000 + 100x + x^2)"

"+y(1400 - 2x -4y)-(10000+6y^2)"

"=900x-2x^2-2xy-7000-100x-x^2"

"+1400y-2xy-4y^2-10000-6y^2"

"=-3x^2-10y^2-4xy+800x+1400y-17000"

b)


"x+y=60"

"\pi(x)=-3x^2-10(60-x)^2-4x(60-x)"

"+800x+1400(60-x)-17000"

"=-9x^2+360x+31000"

"x_{vertex}=-\dfrac{360}{2(-9)}=20"

"y=60-20=40"

The values of "x=20" and "y=40" maximize profit.